OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (4, −7) and a directrix of y = −15. Put the equation in standard form. f(x) = one sixteenthx2 − 8x + 11 f(x) = one sixteenthx2 − 8x − 10 f(x) = one sixteenthx2 − one halfx + 11 f(x) = one sixteenthx2 − one halfx − 10

OpenStudy (anonymous):

f(x)=(1/16)^2 is it written like this?

OpenStudy (anonymous):

Yes, but whats the other half of the equation?

OpenStudy (anonymous):

Not sure yet I was just making sure that's how it was written numerically. Do you know what the focus is? I don't want to just give you the answer, but I'll teach you

OpenStudy (anonymous):

Thanks :) and the focus is (4, -7)

OpenStudy (anonymous):

No I mean do you know the.. hmm definition of the focus? because that is important

OpenStudy (anonymous):

Oh. No, I have this class on flvs and it literally is so difficult to understand, I'm guessing that's the point of the parabola though, like where it starts.

OpenStudy (anonymous):

No, the focus of the parabola is a point inside of it. I'll draw you a diagram. (I took alg. 2 in school, redid it on FLVS and like.. ugh. it sucked. I feel your pain.)

OpenStudy (anonymous):

Thanks :) and I know! It's ridiculous.

OpenStudy (anonymous):

 sorry for that random line

OpenStudy (anonymous):

But the equation for the focus of a vertical parabola is 4p(y-k)=(x-h)^2

OpenStudy (anonymous):

you know what the variables stand for?

OpenStudy (anonymous):

It's ok lol. and ok :) that makes sense and no.

OpenStudy (anonymous):

ok h is the horizontal distance. This might be difficlt without a diagram, let me go find my math book. 10 minutes tops

OpenStudy (anonymous):

Ok tysm

OpenStudy (anonymous):

OMG LIKE WHAT IS LIFE I REMEMBER NOW THIS BOOK I LOVE IT

OpenStudy (anonymous):

sorry okay the directrix is equal y=-p

OpenStudy (anonymous):

here I will copy what the book says

OpenStudy (anonymous):

"You know that the graph of y=ax^2 is a parabola that opens up or down with vertex (0,0) and axis of symmetry is x=0. On any parabola, each point is equidistant from a point called the FOCUS on the line called the DIRECTRIX." does that make sense so far?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

ok the focus lies on the axis of symmetry. the directrix is perpendicular to the axis of symmetry the vertex lies halfway between the focus and directrix.

OpenStudy (anonymous):

Alrighty

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

Alright. The equation the parabola opens up or down ns has vertex of (0,0) can also be written as x^2=4py. (this is where p becomes important). Parabolas also open left and right , where y^2=4px

OpenStudy (anonymous):

the focus and the directrix lie |p| units from the vertex.

OpenStudy (anonymous):